ZombieSlayer:

Check all restrictions on the domain of g circle h. x not-equals 0 x not-equals negative 2 x not-equals negative two-thirds There are no restrictions.

ZombieSlayer:

1 attachment
supie:

What do you think the answer is based on whats given?

ZombieSlayer:

2/3?

ZombieSlayer:

correct?

ZombieSlayer:

or nah lol

jacob28:

its not a

supie:

Yeah.

ZombieSlayer:

is it C. x not-equals negative two-thirds

jimthompson5910:

Hint: \[g(x) = \frac{1}{x+2}\] \[g(h(x)) = \frac{1}{h(x)+2}\] \[g(h(x)) = \frac{1}{3x+2}\] \[(g \circ h) = \frac{1}{3x+2}\]

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ZombieSlayer is it C. x not-equals negative two-thirds \(\color{#0cbb34}{\text{End of Quote}}\) Yes! Correct!

jacob28:

ye

ZombieSlayer:

thank you

supie:

Np.